A cart is propelled over an xy plane with acceleration components ax = 5.7 m/s2 and ay = -2.8 m/s2. Its initial velocity has components v0x = 8.7 m/s and v0y = 12.5 m/s. In unit-vector notation, what is the velocity of the cart when it reaches its greatest y coordinate?
Answer #1

at it's greatest y coordinate,y-component of velocity is zero

So it has only x-component of velocity

Hence


Vy = Voy + (ay*t)

0 = 12.5 -(2.8*t)


t = 4.46 sec is the time taken by the cart to reach greatest y-coordinate


Vx = Vox +(ax*t)

Vx = (8.7)+(5.7*4.46) = 34.12 m/sec